# A Deep Learning-Based Supervised Transfer Learning Framework for DOA Estimation with Array Imperfections

### v2

@@ II-B Signal Model with Array Imperfection
In practical applications, the array signal reception model is often not as ideal as its theoretical counterpart. Position biases may be introduced during the arrangement of array elements. Amplitude and phase inconsistencies may be caused by defects in the design and manufacturing of sensors. And mutual coupling effects may be induced during the signal reception process. In [29], Liu et al. established a simplified model to show how these three array imperfections affect the signal model in ULA, and introduced a hyperparameter $\rho\in[0,1]$ to succinctly control the intensity of these array imperfections.
$\boldsymbol{e}_{\text{pos}}&=\rho\times\left[0,-0.2,\dots,-0.2,0.2,0.2,\dots,0.2\right]^{T}\times d\\ \boldsymbol{e}_{\text{gain}}&=\rho\times\left[0,0.2,\dots,0.2,-0.2,\dots,-0.2\right]^{T}\\ \boldsymbol{e}_{\text{phase}}&=\rho\times\left[0,-30^{\circ},\dots,-30^{\circ},30^{\circ},\dots,30^{\circ}\right]^{T}\\ \boldsymbol{e}_{\text{mc}}&=\rho\times\left[\gamma^{0},\gamma^{1},\dots,\gamma^{M-1}\right]^{T}\end{aligned}$
where $\gamma=0.3e^{j60^{\circ}}$ is the mutual coupling coefficient between adjacent sensors. The term $\boldsymbol{e}_{\text{pos}}$ represents the positional bias of each array element, while $\boldsymbol{e}_{\text{gain}}$ and $\boldsymbol{e}_{\text{phase}}$ denote the amplitude and phase biases, respectively. $\boldsymbol{e}_{\text{mc}}$ models the magnitude of the mutual coupling effect for a ULA, where the strength of the mutual coupling is proportional to the square of the distance. The matrix $\boldsymbol{E}_{\text{mc}}$, denoted as a Toeplitz matrix with parameter vector $\boldsymbol{e}_{\text{mc}}$ [59], can be employed to model the mutual coupling across the entire array:
@@
$\boldsymbol{a}(\theta,\boldsymbol{e})&=\left(\boldsymbol{I}_{M}+\delta_{mc}\boldsymbol{E}_{mc}\right)\times\left(\boldsymbol{I}_{M}+\operatorname{Diag}\left(\delta_{gain}\boldsymbol{e}_{gain}\right)\right)\\ &\times\operatorname{Diag}\left(\exp\left(j\delta_{\text{phase}}\boldsymbol{e}_{\text{phase}}\right)\right)\times\boldsymbol{a}\left(\theta,\boldsymbol{p}+\delta_{pos}\boldsymbol{e}_{pos}\right)\end{aligned}$
where $\delta(\cdot)$ is used to denote specific array imperfections, $I_{M}$ represents the $M\times M$ identity matrix, while $\operatorname{Diag}(\cdot)$ denotes a diagonal matrix formed from a given vector. And $\boldsymbol{a}\left(\theta,\boldsymbol{p}+\boldsymbol{e}_{pos}\right)$ refers to the array steering vector obtained based on the position biases:
@@
By substituting $\boldsymbol{A}(\boldsymbol{\theta},\boldsymbol{e})$ for the steering matrix $\boldsymbol{A}(\boldsymbol{p},\boldsymbol{\theta})$ in Section II-A, a more accurate simulation of the array signal reception model in practical scenarios can be achieved.
@@ III NETWORKS FOR DOA ESTIMATION
Figure 2: The overrall structure of DOA-ViT and the processing procedure of input covariance matrix.
Given the remarkable performance of the Vision Transformer (ViT) model across various domains, its strong nonlinear representation capability and efficient feature learning ability are expected to enhance subsequent transfer learning. Therefore, we adopt ViT as our baseline model. The overall architecture of the ViT model designed for DOA estimation, referred to as DOA-ViT, is illustrated in Figure 2. In terms of model design, we adhere to the original Vision Transformer framework [60] as closely as possible to ensure consistency and leverage its proven architecture.
@@ III-A Covariance Matrix Embedding
Since the transformer encoder can only operate on a set of input vectors, it is necessary to encode the input covariance matrix into a set of vectors.
To achieve this, the split and vectorization operation is employed, which is detailed in (11) and (12). The sampled SCM $\widetilde{\boldsymbol{R}}_{y}$ is partitioned column-wise into $M$ smaller blocks and subsequently each block is flattened into a vector of size $2M$, resulting in $M$ vectors. Finally, a linear layer is utilized to encode each input vector, generating a matrix of dimensions $M\times D$, which represents $M$ vectors, each of dimension $D$.
@@
$\widetilde{\boldsymbol{r}}_{i}$ $=\left[\text{real}(\boldsymbol{r}_{i});\text{imag}(\boldsymbol{r}_{i})\right],\quad\widetilde{\boldsymbol{r}}\in\mathbb{R}^{2M}$ (12)
@@
The formulation of Covariance Matrix Embedding procedure is presented in Eqs. (11) to (14). Notably, similar to the widely used transformer models, a learnable $D$-dimensional vector, called the DOA estimation token, is added to the first column of the embeded matrix. Since the transformer encoder captures global dependencies through self-attention, the DOA estimation token effectively integrates information from the entire sequence and is subsequently used in the final task, such as classification and regression. This approach simplifies model design and enables the transformer model to directly leverage global context for task prediction.
III-B Position Embedding
Transformer models inherently lack information about the order of input vectors. In other words, even if the positions of input vectors are randomly shuffled, the transformer encoder produces identical outputs for each input vector, as it treats the input vectors as an unordered set. To address this limitation, positional embeddings are introduced to encode the relative or absolute positions of the input vectors. Specifically, a set of learnable vectors, known as positional embeddings, is added directly to the input vectors as additional positional encoding information:
@@ III-B Position Embedding
By adding position embeddings to the input representations, the model can incorporate positional information into its computations, ensuring that the order of the input vectors influences the final representation. This mechanism is essential for tasks where understanding the sequential arrangement of input vectors significantly impacts the model’s performance, such as DOA estimation.
III-C Transformer Encoder
The Multi-Head Self-Attention (MHSA) mechanism is a key component of the Transformer Encoder, which is designed to enhance the model’s capacity to capture complex relationships within sequences.
Given an input $\boldsymbol{Z}\in\mathbb{R}^{D\times(M+1)}$, where $D$ is the feature dimension and $M+1$ is the sequence length, the MHSA operates are summarized as follow.
@@ III-C Transformer Encoder
where $W^{O}\in\mathbb{R}^{D\times hd_{k}}$ is the output projection matrix. As illustrated in Fig. 2, the Transformer Encoder, consisting of layer normalization, MHSA and an MLP, enables the DOA-ViT model to effectively capture and integrate global information.
III-D DOA Estimation Head
The DOA Estimation Head serves as the final layer for DOA estimation. In this module, the vector corresponding to the DOA estimation token is extracted; then a linear layer is directly applied to regress the target DOA values. The loss function between the predicted DOA values and the true DOA is expressed as:
@@ III-D DOA Estimation Head
where $\boldsymbol{\psi}$ represents all the parameters of the model and $\hat{\boldsymbol{\theta}}$ is the estimation result of the DOA-ViT model, which includes $K$ DOAs. $\boldsymbol{\theta}$ corresponds to the true DOAs, and the set $\mathcal{P}$ encompasses all $M\times M$ permutation matrices, defined as binary matrices where each row and column contains exactly one non-zero element. The loss function for a dataset $\mathcal{D}$ is described in the following manner:
$\mathcal{L}_{\mathcal{D}}(\boldsymbol{\psi})=\frac{1}{|\mathcal{D}|}\sum_{(\boldsymbol{\theta}^{(j)},\widetilde{\boldsymbol{R}}^{(j)})\in\mathcal{D}}l(\widetilde{\boldsymbol{R}}^{(j)},\boldsymbol{\theta}^{(j)};\boldsymbol{\psi})$ (20)
where $|\mathcal{D}|$ denotes the number of samples in the simulated dataset and $\widetilde{\boldsymbol{R}}^{(j)}$ is the $j$th sampled SCM in the dataset. It is worth noting that our proposed method can be easily extended to arbitrary array configurations and 2D (Two-Dimensional) DOA estimation. If one aims to implement 2D DOA estimation, the DOA Estimation Head can be directly extended to predict the values of $\theta$ and $\phi$. The corresponding loss function for this approach is derived in the following form:
@@ IV Transfer Learning Approach
By leveraging large-scale datasets for training, deep neural networks have achieved superior performance in DOA estimation. However, due to the statistical differences in the sampled SCM between simulated data and real-world data with array imperfections, the model’s performance in practical scenarios is adversely affected.
Inspired by recent advancements in transfer learning [61, 62], and in order to address the significant performance degradation of deep learning models due to array imperfections in practical scenarios, we propose a supervised transfer learning approach, which is designed to enhance the adaptability of deep learning methods to array imperfections.
IV-A Goal of Transfer Learning
Transfer learning typically involves two tasks: the source task and the target task. The source task usually has abundant data available to effectively train the model, whereas the target task has a limited amount of data. Building on machine learning methods, transfer learning allows prior knowledge from the source task to be applied to the target task, even though there are different data distributions between the source and target tasks [55].
IV-B Source and Target Domains Modeling
In our transfer learning approach, two distinct scenarios are considered. The first is the ideal array signal reception scenario, which serves as a scenario that most deep learning models focus on. In this scenario, training data can typically be generated in large quantities through simulation programs, resulting in ample data availability at a relatively low cost. Consequently, deep learning algorithms can be effectively trained and validated based on the extensive dataset, ultimately achieving satisfactory performance.
The second scenario pertains to practical applications, where array imperfections in the antenna system are present. In this case, data should be obtained through actual measurements, leading to higher costs and a relatively limited dataset.
@@ IV-B Source and Target Domains Modeling
The ideal scenario data is treated as source domain data, and the practical scenario data as target domain data. Therefore, how to effectively leverage the prior knowledge acquired from a model trained on a substantial amount of simulation data to enhance model’s performance in a data-scarce or even data-absent practical setting becomes a valuable area for exploration.
IV-C Transfer learning Framework
Assume that we have already trained a model with weight parameters $\boldsymbol{\psi}_{s}$ that performs well on simulated data. This model is capable of effectively extracting features and performing DOA estimation in the source domain. However, due to the presence of array imperfections, substantial statistical discrepancies arise between the data distributions of the source and target domains. Consequently, when applied to the target domain, the model with parameters $\boldsymbol{\psi}_{s}$ extracts features that are significantly influenced by these statistical discrepancies, thereby deteriorating DOA estimation performance.
To mitigate this issue, the objective of this work is to train a feature extractor for the target domain that is designed to learn domain-invariant features, which remain unaffected by domain shifts induced by array imperfections, such as element position biases, amplitude-phase inconsistencies and mutual coupling effects that distort the statistical properties of the sampled SCM. Only the essential information relevant to the DOA estimation task is expected to be retained in domain-invariant features, thereby enhancing the model’s generalization ability in practical scenarios. The framework of our transfer learning framework is illustrated in Fig. II, and we will provide a detailed introduction in the following.
@@ IV-C Transfer learning Framework
In our transfer learning approach, we remove the DOA estimation Head from the DOA-ViT model and treat the remaining structure as a feature extractor, denoted as $\mathcal{F}$, which allows us to extract features from both the source and target domains:
@@
where $\mathcal{F}_{s}$ and $\mathcal{F}_{t}$ represent the feature extractors for the source and target domains, respectively, while $\boldsymbol{\psi}_{s}$ and $\boldsymbol{\psi}_{t}$ represent the model parameters for the source and target domain models. $\boldsymbol{z}_{s,i}$ represents the features extracted from the $i$-th source domain input, and $\boldsymbol{z}_{t,i}$ denotes the features extracted from the $i$-th target domain input.
In practical scenarios, a dataset can be constructed by collecting a series of sampled SCMs along with their corresponding DOA values. Base on the true DOA values, a simulation data generator is employed to generate corresponding ideal covariance matrices that are unaffected by array imperfections, which serve as the source domain data to be aligned. As illustrated in Fig. II, during the transfer learning approach, data in the source domain is fed into $\mathcal{F}_{s}$ with frozen pre-trained weights to ensure the effectiveness of feature extraction from the source domain, while data in the target domain is fed into a trainable feature extraction module $\mathcal{F}_{t}$. In this way, by aligning the features $\boldsymbol{z}_{t,i}$ to $\boldsymbol{z}_{s,i}$, the target domain feature extractor $\mathcal{F}_{t}$ can be guided to extract features that can be directly used for DOA estimation, thereby mitigating the impact of array imperfections. Based on the feature alignment procedure described above, two loss functions are proposed for feature alignment as follows.
MSE Loss: A mean squared error is directly applied between the features of the source and target domains to achieve the alignment of target domain features with source domain features.
@@
[…]
Cosine Similarity Loss: To further align the structure of features between the source and target domains, we treat the features as vectors in a high-dimensional space. A cosine similarity loss is proposed to achieve angular alignment between the source domain vectors and the target domain vectors.
$\mathcal{L}_{\cos}(\boldsymbol{z}_{i}^{S\leftrightarrow T})=\frac{\boldsymbol{z}_{s,i}\cdot\boldsymbol{z}_{t,i}}{\|\boldsymbol{z}_{s,i}\|\cdot\|\boldsymbol{z}_{t,i}\|}$ (24)
where $S$ and $T$ represent the source and target domains, respectively. Finally, the overall loss function can be described as follows:
@@
where $\boldsymbol{Z}^{S}$ and $\boldsymbol{Z}^{T}$ represent all feature vectors extracted from the source domain and target domain, respectively, while $\alpha$ and $\beta$ are hyper-parameters controlling the effect of the two loss functions.
Building upon the loss function of the transfer learning approach, the optimization of the target domain model parameters $\boldsymbol{\psi}_{\text{t}}$ is performed using first-order methods, e.g., Adaptive Moment Estimation (Adam) optimizer. To provide a more precise depiction of the overall process, the transfer learning procedure of DOA-ViT using Adam is detailed as Algorithm 1.
@@
0:  DOA-ViT weights $\boldsymbol{\psi}_{s}$ pretrained on the source domain, hyperparameter $\epsilon$.
@@
1:  Initialize DOA-ViT weights $\boldsymbol{\psi}_{s}$ for source domain model.
2:  Initialize DOA-ViT weights $\boldsymbol{\psi}_{t}$ for target domain model.

### v5

@@ II-B Signal Model with Array Imperfection
In practical applications, the array signal reception model is often deviates from its theoretical counterpart. Such non-ideal conditions typically arise from three primary sources: position biases introduced during array element arrangement; amplitude and phase inconsistencies caused by design and manufacturing defects; and mutual coupling effects induced during the signal reception process. In [35], Liu et al. established a simplified model to demonstrate how these three array imperfections affect the signal model in a ULA, introducing a hyperparameter $\rho\in[0,1]$ to succinctly control the intensity of these effects.
$\begin{aligned} \boldsymbol{e}_{\text{pos}}&=\rho\times\left[0,-0.2,\dots,-0.2,0.2,0.2,\dots,0.2\right]^{T}\times d\\ \boldsymbol{e}_{\text{gain}}&=\rho\times\left[0,0.2,\dots,0.2,-0.2,\dots,-0.2\right]^{T}\\ \boldsymbol{e}_{\text{phase}}&=\rho\times\left[0,-30^{\circ},\dots,-30^{\circ},30^{\circ},\dots,30^{\circ}\right]^{T}\\ \boldsymbol{e}_{\text{mc}}&=\rho\times\left[\gamma^{0},\gamma^{1},\dots,\gamma^{M-1}\right]^{T}\end{aligned}$
where $\gamma=0.3e^{j60^{\circ}}$ is the mutual coupling coefficient between adjacent sensors. The vectors $\boldsymbol{e}_{\text{pos}}$, $\boldsymbol{e}_{\text{gain}}$ and $\boldsymbol{e}_{\text{phase}}$ denote the element positional, amplitude, and phase biases, respectively. Specifically, $\boldsymbol{e}_{\text{mc}}$ models the magnitude of the mutual coupling effect in a ULA, where the strength of the mutual coupling is proportional to the square of the distance between elements. The overall mutual coupling across the entire array is represented by the matrix $\boldsymbol{E}_{\text{mc}}$, which is defined as a Toeplitz matrix parameterized by the vector $\boldsymbol{e}_{\text{mc}}$ [44]:
@@
$\begin{aligned} \boldsymbol{a}(\theta,\boldsymbol{e})&=\left(\boldsymbol{I}_{M}+\delta_{mc}\boldsymbol{E}_{mc}\right)\times\left(\boldsymbol{I}_{M}+\operatorname{Diag}\left(\delta_{gain}\boldsymbol{e}_{gain}\right)\right)\\ &\times\operatorname{Diag}\left(\exp\left(j\delta_{\text{phase}}\boldsymbol{e}_{\text{phase}}\right)\right)\times\boldsymbol{a}\left(\theta,\boldsymbol{p}+\delta_{pos}\boldsymbol{e}_{pos}\right)\end{aligned}$
where $\delta(\cdot)$ indicates the presence of specific array imperfections. $\boldsymbol{I}_{M}$ represents the $M\times M$ identity matrix, while $\operatorname{Diag}(\cdot)$ denotes a diagonal matrix formed from a given vector. And $\boldsymbol{a}\left(\theta,\boldsymbol{p}+\boldsymbol{e}_{pos}\right)$ refers to the array steering vector incorporating the position biases:
@@
By substituting $\boldsymbol{A}(\boldsymbol{\theta},\boldsymbol{e})$ for the steering matrix $\boldsymbol{A}(\boldsymbol{p},\boldsymbol{\theta})$ introduced in Section II-A, a more accurate simulation of the array signal reception model in practical scenarios can be achieved.
@@ III NETWORKS FOR DOA ESTIMATION
Figure 3: The overall structure of TansDOA and the processing procedure of input covariance matrix.
To capture global dependencies among array elements, we propose a Transformer-based model named TransDOA. The overall framework comprising the TransDOA architecture and the transfer learning strategy is illustrated in Fig. 2. Specifically, the detailed structure of the TransDOA is further presented in Fig. 3.
@@ III-A Covariance Matrix Embedding
To accommodate the Transformer’s sequence-based architecture for DOA estimation, the SCM $\widetilde{\boldsymbol{R}}_{y}$ is partitioned and mapped into a sequence of vectors. Specifically, $\widetilde{\boldsymbol{R}}_{y}$ is split column-wise into $M$ blocks, each flattened and transformed into a $D$-dimensional embedding $\boldsymbol{x}_{i}$ via a linear layer $\boldsymbol{E}$, as formulated in (10)–(12). A learnable DOA token is prepended to the input sequence to aggregate global features for the final estimation task (13).
@@
$\widetilde{\boldsymbol{r}}_{i}$ $=\left[\text{real}(\boldsymbol{r}_{i});\text{imag}(\boldsymbol{r}_{i})\right],\quad\widetilde{\boldsymbol{r}}_{i}\in\mathbb{R}^{2M}$ (11)
@@
Since the self-attention mechanism is permutation-invariant, learnable positional embeddings $\boldsymbol{E}_{pos}$ are added to the input representations to preserve the spatial structure of the array signal:
@@
Further details regarding the SCM Embedding process and the comprehensive configuration of the TransDOA model are presented in Fig. 3.
III-B Transformer Encoder
The Multi-Head Self-Attention (MHSA) mechanism serves as the core component of the Transformer Encoder, designed to capture complex global dependencies within the input sequence.
Given an input $\boldsymbol{Z}\in\mathbb{R}^{D\times(M+1)}$, where $D$ is the feature dimension and $M+1$ is the sequence length, the MHSA operates are defined as follow.
@@ III-B Transformer Encoder
where $W^{O}\in\mathbb{R}^{D\times hd_{k}}$ is the output projection matrix. As illustrated in Fig. 3, the Transformer Encoder, consisting of layer normalization, MHSA and an MLP, enables the TransDOA model to effectively capture and integrate global information.
III-C DOA Estimation Head
The DOA Estimation Head serves as the final regression module. The feature vector corresponding to the aggregated DOA token is extracted and projected via a linear layer to predict the target DOAs. To address the permutation ambiguity inherent in multi-source estimation, we employ the permutation invariant training loss:
@@ III-C DOA Estimation Head
where $\boldsymbol{\psi}$ denotes the model parameters and $\hat{\boldsymbol{\theta}}$ represents the estimation result, which includes $K$ DOAs. $\boldsymbol{\theta}$ corresponds to the true DOAs, and the set $\mathcal{P}$ encompasses all $K\times K$ permutation matrices, defined as binary matrices where each row and column contains exactly one non-zero element.
Furthermore, the proposed framework extends naturally to 2D (Two-Dimensional) DOA estimation. By extending the DOA Estimation Head to predict both azimuth $\theta$ and elevation $\phi$ angles, the joint loss function is formulated as:
@@ IV Transfer Learning Approach
Deep learning models trained on large-scale simulated datasets often suffer from performance degradation in practical scenarios. This is primarily due to the domain gap caused by array imperfections.
To bridge this gap, we propose a supervised transfer learning framework based on feature alignment. As illustrated in Fig. III, our approach aims to train a target domain feature extractor that maps distorted SCM into a domain-invariant feature space aligned with the ideal array manifold.
IV-A Domain Characterization
In our approach, two distinct scenarios are considered. The first is the ideal reception scenario, a standard paradigm in deep learning research. In this scenario, large-scale training datasets can be generated via simulation at low cost, enabling effective model training and superior performance.
The second scenario pertains to practical applications characterized by array imperfections. In contrast to the ideal case, data acquisition here relies on physical measurements, inevitably resulting in high costs and data scarcity. Based on these characteristics, we formally define the respective domains as follows:
  • Source Domain ($\mathcal{D}_{s}$): Consists of abundant ideal SCMs generated via simulation. A model trained here captures the theoretical array manifold but fails on imperfect data.
  • Target Domain ($\mathcal{D}_{t}$): Consists of limited labeled data from the practical system with array imperfections.
@@ IV-A Domain Characterization
How to effectively leverage the prior knowledge acquired from a model trained on a substantial amount of simulation data to enhance model’s performance in a data-scarce or even data-absent practical setting becomes a valuable area for exploration.
IV-B Transfer learning Framework
To address the performance degradation caused by statistical discrepancies between the source and target domains, we propose a transfer learning framework, as illustrated in Fig. III. By leveraging the pre-trained source parameters $\boldsymbol{\psi}_{s}$, our objective is to train a target extractor that learns domain-invariant features, thereby ensuring robustness against array imperfections.
@@ IV-B Transfer learning Framework
Specifically, we remove the DOA estimation Head from the TransDOA model and utilize the backbone as a feature extractor, denoted as $\mathcal{F}$. The feature representations for the source and target domains are obtained as:
@@
where $\mathcal{F}_{s}$ and $\mathcal{F}_{t}$ are the feature extractors for the source and target domains, parameterized by $\boldsymbol{\psi}_{s}$ and $\boldsymbol{\psi}_{t}$, respectively. The vectors $\boldsymbol{z}_{s,i}$ and $\boldsymbol{z}_{t,i}$ represent the feature extracted from the $i$-th pair of source and target inputs.
To ensure exact feature correspondence between the source and target domains, it is critical to construct a strictly paired dataset. A Simulation Data Generator is employed for this purpose. In practical scenarios, we collect a series of SCMs along with their corresponding DOA values. Subsequently, these true DOA values are fed into the Simulation Data Generator to generate corresponding ideal covariance matrices. This process establishes a one-to-one mapping, where each collected real-world sample is paired with a generated ideal counterpart, thereby forming the requisite paired dataset for effective knowledge transfer and model calibration.
Utilizing the constructed paired dataset, the transfer learning framework is illustrated in Fig. III. The pre-trained source feature extractor $\mathcal{F}_{s}$ is frozen to serve as a stable reference. The target encoder $\mathcal{F}_{t}$ is then trained to align its extracted features $\boldsymbol{z}_{t,i}$ with the reference features $\boldsymbol{z}_{s,i}$. This alignment compels $\mathcal{F}_{t}$ to suppress domain-specific distortions caused by array imperfections. To enforce such robust alignment, we propose a composite objective function consisting of two terms:
1) MSE Loss: The Mean Squared Error (MSE) is employed to minimize the Euclidean distance between the feature representations of the two domains:
@@
where $S$ and $T$ represent the source and target domains, respectively. The notation $\|\cdot\|$ denotes the $l_{2}$-norm of a vector.
2) Cosine Similarity Loss: To further enforce structural consistency, we treat features as vectors in a high-dimensional space and align their geometric directions. The cosine loss is formulated as:
$\mathcal{L}_{\cos}(\boldsymbol{z}_{i}^{S\leftrightarrow T})=1-\frac{\boldsymbol{z}_{s,i}\cdot\boldsymbol{z}_{t,i}}{\|\boldsymbol{z}_{s,i}\|\cdot\|\boldsymbol{z}_{t,i}\|}$ (22)
By minimizing this loss, we maximize the cosine similarity, ensuring the target features share the same angular direction as the source features. Finally, the overall objective function is a weighted sum of the aforementioned losses:
@@
where $\boldsymbol{Z}^{S}$ and $\boldsymbol{Z}^{T}$ represent all feature vectors extracted from the source domain and target domain, respectively, while $\alpha$ and $\beta$ are hyper-parameters controlling the effect of the two loss functions. To provide a more precise depiction of the overall process, the transfer learning procedure of TransDOA using Adam is detailed as Algorithm 1.
@@
0: TransDOA weights $\boldsymbol{\psi}_{s}$ pretrained on the source domain, hyperparameter $\epsilon$.
@@
1: Initialize TransDOA weights $\boldsymbol{\psi}_{s}$ for source domain model.
2: Initialize TransDOA weights $\boldsymbol{\psi}_{t}$ for target domain model.